### v1

@@ 2.2 Thermal, Fast and Interactive Statistical Toolkit (Thermal-FIST)
For the present investigation the grand canonical ensemble (GCE), including light nuclei in the particle list, is used to fit the data, while strangeness suppression is modeled using a standard $\gamma_{s}$ suppression parameter. The conditions of (average) strangeness neutrality and a total charge to baryon ration of $Q/A=0.4$ is enforced by introducing the strange and electric-charge chemical potentials $\mu_{S}$ and $\mu_{Q}$ as fit parameters.  [^2] To check, we have also applied the strangeness canonical version of the model and found that it provides very similar results. To perform $\chi^{2}$ fits one also needs to provide (statistical) error bars. While the statistics in the simulations can in principle be increased arbitrarily, minimizing the statistical error, here we nevertheless assume a $10\%$ statistical error for all hadronic species, similar to the typical systematic error values found in the experiment. This avoids introducing a bias to extremely rare particle species.
@@ 4 Summary
We have studied the dependence of thermal fit parameters on the choice of hadrons used as input for the fit as well as the underlying equation of state. We have found systematic effects from both which need to be taken into account when one wants to relate observables at certain beam energies to features of the QCD phase diagram using the thermal freeze-out line as reference. Inconsistencies in the underlying equation of state may lead to an uncertainty of $\approx 10$ MeV in the extracted temperature and $\approx 50$ MeV in chemical potential. The effects from an incomplete set of hadrons in the fit may even be larger. In addition, if a different equation of state is used for the fitting as for the evolution of the system, a shift in the beam energy at which a certain chemical freeze-out point is reached can be shifted by approximately 0.5 GeV which corresponds to more than $10\%$ of the beam energy in the center of mass frame. These systematic uncertainties need to be taken seriously if one wants to locate important features of the phase diagram from heavy ion collisions.
@@ Acknowledgments
T.R. gratefully acknowledges financial support by the Fulbright U.S. Scholar Program, which is sponsored by the U.S. Department of State and the German-American Fulbright Commission. This article’s contents are solely the responsibility of the authors and do not necessarily represent the official views of the Fulbright Program, the Government of the United States, or the German-American Fulbright Commission. T.R. gratefully acknowledges support from The Branco Weiss Fellowship - Society in Science, administered by the ETH Zürich. V.V. has been supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, Early Career Research Program under Award Number DE-SC0026065. The computational resources for this project were provided by the Center for Scientific Computing of the GU Frankfurt and the Goethe–HLR and GSI green cube.

### v2

@@ 2.2 Thermal, Fast and Interactive Statistical Toolkit (Thermal-FIST)
For the present investigation the grand canonical ensemble (GCE), including light nuclei in the particle list, is used to fit the data, while strangeness suppression is modeled using a standard $\gamma_{s}$ suppression parameter. The conditions of (average) strangeness neutrality and a total charge to baryon ration of $Q/A=0.4$ is enforced by introducing the strange and electric-charge chemical potentials $\mu_{S}$ and $\mu_{Q}$ as fit parameters.  [^2] To check, we have also applied the strangeness canonical version of the model and found that it provides very similar results. To perform $\chi^{2}$ fits one also needs to provide (statistical) error bars. While the statistics in the simulations can in principle be increased arbitrarily, minimizing the statistical error, here we nevertheless assume a $10\%$ statistical error for all hadronic species, similar to the typical systematic error values found in the experiment. This avoids introducing a bias to extremely rare particle species. For all the fits in the following we always followed the ’standard’ procedure and employed the HRG model without any modifications to the EoS.
@@ 3 Results
Even though the underlying equation of state used in the simulation with the CMF equation of state (leading to changes in the particle yields) is not consistent with the equation of state used for the fitting, a satisfactory fit can still be obtained. Similarly, a HRG fit to real data may give a good fit result but not the ’correct’ T-$\mu_{B}$-values.
@@ 4 Summary
We have studied the dependence of thermal fit parameters on the choice of hadrons used as input for the fit as well as the underlying equation of state. We have found systematic effects from both which need to be taken into account when one wants to relate observables at certain beam energies to features of the QCD phase diagram using the thermal freeze-out line as reference. Changes in the underlying equation of state may lead to a shift of $\approx 10$ MeV in the extracted temperature and $\approx 50$ MeV in the chemical potential. Even with an inconsistency in EoS between the thermal fit and underlying simulation (or measured data), a satisfactory fit can be obtained, albeit with systematically shifted freeze-out parameters. This is not a trivial result as it is not self-evident that such a good fit should still be possible with a change in the equation of state or that this would lead to the same freeze-out curve. It also means that there is no unique (EoS-independent) relation between a beam energy and the expected freeze-out points. The effects from an incomplete set of hadrons in the fit may even be larger. In addition, if a different equation of state is used for the fitting as for the evolution of the system, a shift in the beam energy at which a certain chemical freeze-out point is reached can be shifted by approximately 0.5 GeV which corresponds to more than $10\%$ of the beam energy in the center of mass frame. These systematic uncertainties need to be taken seriously if one wants to locate important features of the phase diagram from heavy ion collisions.
@@ Acknowledgments
T.R. gratefully acknowledges financial support by the Fulbright U.S. Scholar Program, which is sponsored by the U.S. Department of State and the German-American Fulbright Commission. This article’s contents are solely the responsibility of the authors and do not necessarily represent the official views of the Fulbright Program, the Government of the United States, or the German-American Fulbright Commission. T.R. gratefully acknowledges support from The Branco Weiss Fellowship - Society in Science, administered by the ETH Zürich. V.V. has been supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, Early Career Research Program under Award Number DE-SC0026065. The computational resources for this project were provided by the Center for Scientific Computing of the GU Frankfurt and the Goethe–HLR and GSI green cube. The publication is funded by the Open Access Publishing Fund of GSI Helmholtzzentrum fuer Schwerionenforschung.